### v1

# Galley: Modern Query Optimization for Sparse Tensor Programs

Abstract.
The tensor programming abstraction has become the key . This framework allows users to write high performance programs for bulk computation via a high-level imperative interface. Recent work has extended this paradigm to sparse tensors (i.e. tensors where most entries are not explicitly represented) with the use of sparse tensor compilers. These systems excel at producing efficient code for computation over sparse tensors, which may be stored in a wide variety of formats. However, they require the user to manually choose the order of operations and the data formats at every step. Unfortunately, these decisions are both highly impactful and complicated, requiring significant effort to manually optimize. In this work, we present Galley, a system for declarative sparse tensor programming. Galley performs cost-based optimization to lower these programs to a logical plan then to a physical plan. It then leverages sparse tensor compilers to execute the physical plan efficiently. We show that Galley achieves high performance on a wide variety of problems including machine learning algorithms, subgraph counting, and iterative graph algorithms.
Query Optimization, Sparse Tensors, Array Programming, Program Optimization
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Berlin, Germanyisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query optimizationccs: Software and its engineering Domain specific languagesccs: Mathematics of computing Mathematical software
1. Introduction
In recent years, the tensor programming model has become ubiquitous across different fields of computation. Popularized by its use in deep learning, it has been applied to problems such as relational query processing (He et al., 2022; Asada et al., 2022; Koutsoukos et al., 2021), data cleaning (Sun et al., 2023), graph algorithms (Szárnyas et al., 2021), and scientific computing (Wang et al., 2022; Stonebraker et al., 2013; Modi et al., 2021) among others. This model promises a high-level imperative abstraction which is highly efficient as long as the problem can be posed in terms of tensor (i.e. array) operations. By doing so, it insulates the user from the many low level concerns that are crucial to achieving good performance. This has allowed experts in fields like data science and machine learning to take advantage of a wide range of hardware infrastructure without having to become experts in high performance computing as well.
Traditional tensor processing frameworks are built on collections of hand-optimized functions, called kernels, which each compute an operation over dense tensors (Abadi et al., 2016; Paszke et al., 2019; Harris et al., 2020; Anderson et al., 1990). The operation can be simple like a matrix-vector multiplication, or it can be a fusion of multiple semantic operations, like $A(B+C)$. However, most data is fundamentally sparse (i.e. most entries are a fill value like 00) such as graph data, one-hot encodings, relational data, meshes in physical simulations, sparse neural networks, and others. Even materializing these datasets as dense arrays can be prohibitively costly, so it’s crucial to perform the computation over its compressed, sparse format (Kjolstad et al., 2017).
To support sparse data, each framework needs to offer many distinct implementations of each operation, one for each combination of input formats. Each input can either be stored densely or in one of many sparse formats, e.g. CSR, CSC, COO. This has lead to an explosion of required kernels as the number of operations and formats both continue to grow. Understandably, these frameworks have been unable to keep up with this implementation effort, resulting in spotty coverage for operations over sparse data (Ivanov et al., 2023).
Example 0.
Consider logistic regression over $n$ data points and $d$ features where $X\in\mathbb{R}^{n\times d}$ is the feature matrix and $\theta\in\mathbb{R}^{d}$ is the parameter vector. Inference is defined as
(1) $P_{i}=\sigma(\sum_{j}X_{ij}\theta_{j})$
where $\sigma$ is the sigmoid function. To compute $Y_{i}=\sum_{j}X_{ij}\theta_{j}$ in the dense case, one could use the ’matmul’ function from the Numpy library. On these inputs, this function specializes to an efficient, hand-coded implementation of dense matrix-vector multiplication. However, if any combination of $X$, $\theta$, and $Y$ can be sparse, Numpy needs to provide eight distinct implementations of matrix-vector multiplication.
To address this issue, the scientific computing community has adopted sparse tensor compilers (STCs) (Kjolstad et al., 2017; Bik et al., 2022; Ahrens et al., 2024; Shaikhha et al., 2022; Hu et al., 2019). These compilers take in a high level imperative tensor program and, separately, a description of the input tensors’ formats [^1] Some systems separate declarative and imperative concerns with a scheduling language. However, the user still controls both aspects.. They automatically produce an efficient kernel implementation in low level code (LLVM, MLIR, etc.). Thus, STCs offer a form of data independence, by allowing the user to separately specify the algorithm and the data layout. For example, Fig. 1 shows a kernel definition for Ex. 1.1 written in Finch, an STC language (Ahrens et al., 2024). The Dense/Sparse input formats are specified in line 1 and are separate from the imperative code in lines 5-15.
However, STCs require users to make a series of complex decisions to produce an efficient implementation. First, the user needs to break their program into a sequence of aggregations. Suppose that the feature matrix in Ex. 1.1 is the result of a matrix multiplication, $X=A\cdot B$, and Eq. (1) becomes $Y_{i}=\sigma(\sum_{jk}A_{il}B_{lj}\theta_{j})$. The user must choose to sum over $j$ or $l$ first, resulting in vastly different runtimes (Ahrens et al., 2022; Zhang et al., 2021). Then, for each aggregate, the user needs to choose the output format, loop order, and access protocols. Consider Fig 1. The user must choose the output format for the intermediate R (line 3). In this case, she chose a Dense format rather than a Sparse format which would be 10$\approx$ 10$\times$ 10 × slower. Then, the user chooses the loop order (lines 5-6). She chose $i$-then-$j$ which is asymptotically faster than $j$-then-$i$ because each out-of-order access to $X$ requires a full scan of the tensor. Lastly, she picks a merge algorithm for each loop that describes how to iterate through the non-zero indices (line 8). Here, $X$ is iterated through and each non-zero $j$ is looked up in $\theta$. If she chose to iterate through $\theta$, each inner loop would scan the entire vector. While STCs allow users to separate the algorithm from the data formats, the user is still responsible for optimizing her program to achieve good performance.
 0. # Specified format for input tensors
 1. FUNC log_regression(X::Dense(Sparse()), $\theta$::Dense())
 2.    # Manually defined intermediate format
 3.    R = Dense()
 4.    # Manually defined loop order
 5.    FOR i=_
 6.        FOR j=_
 7.            # Manually defined access protocols
 8.            R[i] += X[i::iter,j::iter]*$\theta$[j::lookup]
 9.        END
10.    END
11.    P = Dense()
12.    FOR i=_
13.        P[i] = $\sigma$(R[i::iter])
14.    END
15. END
Figure 1. Logistic Regression Impl. in Simplified Sparse Tensor Compiler Language.
In this paper, we propose Galley, a system for declarative sparse tensor programming. Users write tensor programs in a declarative language based on the Einsum notation, similar to Equation (1), and Galley automatically produces an optimized STC implementation. First, it restructures the program into a sequence of aggregation steps, minimizing the total computation and materialization costs (Sec. 5). Then, Galley optimizes each step by selecting the optimal formats for all intermediate tensors, the loop order, and the access method for each index access (Sec. 6). This is all guided by a system for estimating sparsity via statistics on the input tensors (Sec. 7). Galley builds on fundamental principles from cost-based query optimization while developing new techniques that are specific to producing optimized code for sparse tensor compilers(Khamis et al., 2016).
In the first phase, Galley rewrites the input program into a sequence of aggregation steps by adapting the variable elimination framework. This reduces a complex rewriting problem to the problem of finding an optimal elimination order. However, this approach has previously only been defined for sum-product queries. To handle arbitrary sparse tensor algebra programs, we extend it to non-distributive functions (e.g. $\sum_{i}max(A_{ij},B_{j})$), disjunctive functions (e.g. $\sum_{i}A_{i}+B_{i}=\sum_{i}A_{i}+\sum_{i}B_{i}$), and internal aggregates (e.g. $\sum_{j}A_{j}\sqrt{\sum_{k}B_{jk}}$). In Sec. 9, we show that these optimizations produce up to a $\mathbf{100\times}$ speedup for machine learning algorithms over composite feature matrices. After this phase, our program has been converted to a sequence of aggregates over pointwise expressions, i.e. the Logical Plan dialect in Fig. 4.
Next, Galley’s physical optimizer transforms each aggregate into an efficient STC kernel by choosing the loop order, output format, and merge algorithm. Because sparse outer loops avoid inner iterations, a good loop ordering can significantly improve performance. Galley finds the optimal loop order with a novel combination of branch-and-bound and dynamic programming techniques. Following the modern tensor format abstraction (i.e. the fibertree abstraction), Galley then chooses a format for each dimension of the output by examining both the sparsity and write pattern (random vs sequential)(Sze et al., 2017; Chou et al., 2018; Bik et al., 2022). Lastly, for each loop index, we select one input to iterate over and perform lookups on the others.
All of Galley’s optimizations are guided by a framework for estimating the sparsity of intermediate results from statistics on the inputs. The core of this is a minimal interface for composing statistics about inputs to infer statistics about outputs. This makes incorporating a new kind of sparsity statistic as simple as implementing 5 functions to; 1) collect statistics on inputs 2-3) handle conjunctive and disjunctive point-wise operations 4) handle aggregates and 5) produce sparsity estimates. We demonstrate this by implementing a naive estimator based on uniformity assumptions and a worst-case estimator based on cardinality bounds.
Example 0.
In Example 1.1, we assumed that $X$ is given directly. However, feature matrices are generally assembled from more basic inputs. Suppose $S_{ipc}\in\mathbb{B}^{n_{i}\times n_{p}\times n_{c}}$ is a boolean sparse tensor where each non-zero entry is a sale, $i$, of product, $p$, to customer, $c$. $P_{pj}\in\mathbb{R}^{n_{p}\times d}$ and $C_{cj}\in\mathbb{R}^{n_{c}\times d}$ are smaller matrices holding numeric features about products and customers. The feature dimension, indexed by $j$, represents both product and customer features, but each feature comes from either $P_{pj}$ or $C_{cj}$. So, each column is non-zero in either $P$ or $C$, and they are concatenated by addition. $X$ is now defined as $X_{ij}=\sum_{pc}S_{ipc}(P_{pj}+C_{cj})$. Galley’s logical optimizer can take advantage of $X$’s structure by pushing $\theta$ down into its definition,
$Y_{i}=\sigma(\sum_{jpc}S_{ipc}P_{pj}\theta_{j}+\sum_{jpc}S_{ipc}C_{cj}\theta_{j})$
We can then push down the summations,
(2) $Y_{i}=\sigma\left(\sum_{pc}\left(S_{ipc}\sum_{j}\left(P_{pj}\theta_{j}\right)\right)+\sum_{pc}\left(S_{ipc}\sum_{j}\left(C_{cj}\theta_{j}\right)\right)\right)$
By doing this, Galley only materializes vector intermediates rather than the full feature matrix. On similar examples, Sec. 9 shows that this can result in up to 100×100$\times$100 × faster execution.
After logical optimization, Galley’s physical optimizer produces a single STC kernel for each aggregate. For $\sum_{j}P_{pj}\theta_{j}$, it would produce,
    I1 = Dense()
    FOR p=_
        FOR j=_
            I1[p] += P[p::iter,j::iter]*$\theta$[j::lookup]
        END
    END
Contributions We claim the following contributions in this work,
  • We present Galley, a system for declarative sparse tensor programming (Sec.4). Galley automatically translates high-level sparse tensor programs to efficient imperative programs targeting a sparse tensor compiler.
  • Galley supports a highly expressive language for sparse tensor algebra with arbitrary algebraic operators, aggregates within expressions, and multiple outputs.
  • Galley is the first system to perform cost-based logical optimization, redistributing aggregates and reorganizing the pointwise statements in the program. (Sec.5)
  • Galley is the first system to perform cost-based physical optimization to determine loop orders, tensor formats, and merge algorithms for each dimension (Sec.6)
  • We propose a minimal interface for sparsity estimation and demonstrate it by implementing two estimators (Sec.7)
  • We evaluate Galley on several workloads and show that it is up to 100x faster than hand-optimized kernels for mixed dense-sparse workloads and up to 100x faster than a state of the art database for highly sparse workloads.
Figure 2. Fiber Tree Format Abstraction
2. Related Work
Our work differs from other proposals on cost-based optimization for tensor processing due to its targeting of STCs and its expressive input language. SystemDS, formerly SystemML, is a system for end-to-end machine learning over matrices and tabular data (Boehm et al., 2016, 2020; Baunsgaard and Boehm, 2023). It primarily takes in linear algebra (LA) programs and targets a combination of LA libraries and distributed computing via Spark. Later work, SPORES, extended its logical optimizer to leverage relational algebra when optimizing sum-product expressions(Wang et al., 2020). Their core insight was that LA rewrites, which always match and produce 0-2D expressions, are not sufficient. Optimal rewrites need to pass through higher order intermediate expressions. Other work translated sum-product expressions to SQL to leverage highly efficient database execution engines(Blacher et al., 2023). These systems can perform well for highly sparse inputs but struggle on mixed dense-sparse workloads. Tensor relational algebra proposes a relational layer on top of dense tensor algebra which provides a strong foundation for automatically optimizing distributed dense tensor computations (Yuan et al., 2021). In the compilers community, there have been attempts to automatically optimize sparse tensor kernels based on asymptotic performance analyses(Ahrens et al., 2022; Dias et al., 2023). These systems each target a different execution context and focus on different aspects of optimization. Galley expands on this line of work by targeting a new execution engine, proposing novel optimization techniques, and handling a wider range of tensor programs.
3. Background
3.1. Tensor Index Notation
The input to our system, Galley, is an extended version of Einsum notation which we call tensor index notation. Traditional Einsum notation allows for a single summation wrapped around a multiplication operation. For instance, you can describe triangle counting in a graph where $E_{ij}$ is the adjacency matrix with the following Einsum statement,
$t=\sum_{ijk}E_{ij}E_{jk}E_{ik}$
To capture the diverse workloads of tensor programming, we additionally allow the use of arbitrary functions for both aggregates and pointwise operations, nesting of aggregates and pointwise operations, and defining multiple outputs. For example, a user could perform logistic regression to predict entities that might be laundering money. Then, they could filter this set based on whether they occur in a triangle in the transactions graph. This is represented by $\max_{jk}(E_{ij}E_{jk}E_{ik})$ which is $1$ if $i$ occurs in at least one triangle and 00 otherwise.
[…]
$V_{i}$ $=L_{i}\cdot\max_{jk}(E_{ij}E_{jk}E_{ik})$
Tensor compilers like Halide, TACO, and Finch each build off of similar core notations, adding additional structures like FOR-loops to allow users to specify algorithmic choices(Ragan-Kelley et al., 2013; Kjolstad et al., 2017; Ahrens et al., 2024). Crucially, the vast majority of operations in array programming frameworks like Numpy can be expressed as operations in tensor index notation So, while we focus on this notation in this paper, traditional tensor workflows can be captured and optimized in this framework.
Figure 3. Galley Overview
3.2. Sparse Tensor Compilers
In the last decade, the compilers community has developed a series of sparse tensor compilers and shown that they can produce highly efficient code for sparse tensor computations. We use this work as our execution engine, so we take a moment here to explain their important concepts.
Tensor Formats
There are many different ways to represent sparse tensors, and the best approach depends on the data distribution and the workload. Work in this space has converged on the fiber tree abstraction for describing the space of formats(Kjolstad et al., 2017). In this formalism, a tensor format is a nested data structure like the one in Fig. 2. Each layer stores the non-fill (e.g. non-zero) indices in a particular dimension, conditioned on the earlier dimensions, and pointers to the next dimension’s non-fill indices. These layers can be represented in any format that allows for iteration and lookup. In this work, we consider sorted lists, hash tables, bytemaps, and dense vectors, and they each have different performance characteristics in terms of iteration, lookup, and memory footprint. For example, compressed sparse row (CSR) is a common format for sparse matrices. It stores the row dimension as dense vector where each entry points to the set of non-zero columns for that particular row. This set of non-zero columns is then stored in a sorted list, i.e. in a compressed sparse format. Importantly, this abstraction requires tensors to be accessed in the order that they are stored (e.g. row-then-column in the case of CSR) which restricts the set of valid loop orders as we describe next.
A.   Input Program
Plan := Query...    Query := (Name, Expr)
Agg := (Op, Idx..., Expr)   Map := (Op, Expr...)
Expr := Agg | Map | Input | Alias
Input := Tensor[Idx...]  Alias := Name[Idx...]
B.   Logical Plan
Plan := Query...   Query := (Name, Agg)
Agg := (Op, Idx..., Expr)   Map := (Op, Expr...)
Expr := Map | Input | Alias
Input := Tensor[Idx...]  Alias := Name[Idx...]
C.   Physical Plan
Plan := Query...   Query := (Name, Mat, Idx...)
Mat := (Format..., Idx..., Agg)
Agg := (Op, Idx..., Expr)   Map := (Op, Expr...)
Expr := Map | Input | Alias
Input := Tns[PIdx...]  Alias := Name[PIdx...]
PIdx := Idx::Protocol
Figure 4. Query Plan Dialects
Loop Execution Model
The input to a Sparse Tensor Compiler is a high-level domain specific language (DSL) which consists of for-loops, in-place aggregates (e.g. +=+=+ =), and arithmetic over indexed tensors (e.g. $A[i,j]*B[j,k]$). Crucially, the for-loops in these expressions are not executed in a dense manner. Instead, these compilers analyze the input formats and the algebraic properties of the expression to determine which index combinations will produce non-fill entries. In Fig. 1, because 00 is the annihilator of multiplication (i.e. $x*0=0$), only the values of $i$ which map to non-zero entries in $X$ and $\theta$ need to be processed. All other index values will return a zero. So, the outer loop is compiled to an iteration over the intersection of the non-zero $i$ indices in $X$ and $\theta$. Considering the illustration in Fig. 2, we can see how this is simply co-iteration over the top levels of their formats. The inner loop then iterates over the $j$ indices which are non-zero in $X[i,\_]$, i.e. the non-zero columns that occur in each row.
Merge Algorithms
Once the compiler has determined which tensors’ non-zero indices need to be merged to iterate over a particular index, there are several algorithms it can apply. All formats allow for both ordered iteration and lookup operations, so one algorithm is to iterate through the indices of all inputs, similar to a merge join. This is highly efficient per operation. However, it’s linear in the total size of all inputs, even if one is much smaller than the others. Another method is to iterate through one input’s level and lookup that index in the others. In this work, we generally take the latter approach, as described in Sec. 6.3. We refer to the mode of an individual tensor (such as “iterate” or “lookup”) as an access protocol and the overall strategy as a merge algorithm(Ahrens et al., 2023).
4. Galley Overview
In this section, we provide a birds-eye view of Galley and show how an input program is transformed, first to a logical plan then to a physical plan then to an STC program, as illustrated in Fig.  3. The first transformation from input program to logical plan is handled by Galley’s logical optimizer, which extends the variable elimination framework to break the program down into a sequence of aggregations. The second transformation, from logical plan to physical plan, is performed by Galley’s physical optimizer, which determines the loop order, tensor formats, and merge algorithms for each logical query. These steps are each represented by a dialect of our query plan language whose grammar is defined in Fig. 4. In the following discussion, we use this grammar as a guide and show how our example program, logistic regression, would be transformed through these steps.
The input program dialect is equivalent to the tensor index notation defined in Sec. 3.1. Pointwise functions like $A_{ij}*B_{jk}$ are represented with Map. Aggregates like $\sum_{i}$ are denoted by Agg. Each assignment is a Query, and previous assignments are referenced with an Alias. Our logistic regression example from Eq. (1) is defined in this dialect as follows,
Query(P, Map($\sigma$, Agg(+, j, Map(*, X[i,j], $\theta$[j]))))
Note that this notation is compatible with array APIs like Numpy which don’t have named indices. Operations like ’matmul’ can be automatically mapped into this language by generating index names for inputs on the fly and renaming whenever operations imply equality between indices. Also, note that aggregates can be over multiple indices, e.g. Agg(+, i, j, k, ...).
4.1. Logical Plan
The first task in our optimization pipeline, handled by the logical optimizer, is to break down the input program into a sequence of simple aggregates. This is enforced by converting the input program to a logical plan. This dialect is a restriction of the input dialect where each query contains a single aggregate statement that wraps an arbitrary combination of Map, Input, and Alias statements. Intuitively, each logical query corresponds to a single STC kernel that produces a single intermediate tensor, but it does not specify details like loop orders and output formats. To perform this conversion soundly, each input query must correspond to a logical query which produces a semantically equivalent output. To do this efficiently, it must minimize the total cost of all queries in the logical plan.
As an example, our logistic regression program above can be translated into the following logical plan,
Query(R, Agg(+, j, Map(*, X[i,j], $\theta$[j])))
Query(P, Agg(no-op, Map($\sigma$, R[i])))
In this plan, the first query isolates the sum over the $j$ index while the second query performs the remaining sigmoid operation on the result. Note that the latter query uses a no-op aggregate to represent an element-wise operation while conforming to the logical dialect.
4.2. Physical Plan
Given the logical plan, Galley’s physical optimizer determines the implementation details needed to convert each logical query to a STC kernel. Specifically, it defines the loop order of each compiled kernel, the format of each output, and the merge algorithm for each index. As above, this is expressed by converting the logical plan to a physical plan described in the most verbose and constrained dialect. To avoid out-of-order accesses, we require that the index order of inputs and aliases are concordant with the loop order, so the physical optimizer may insert additional queries to transpose inputs. Therefore, each logical query corresponds to one or more physical queries.
Using this language, we can precisely express the program from Fig. 1 as follows where it means iterate and lu means lookup.
Query(R,Mat(dense,i,Agg(+,j,Map(*, X[i::it,j::it],
                                            $\theta$[j::lu]))), i, j)
Query(P,Mat(dense,i, Map($\sigma$, P1[i::it])), i)
The first query computes the sum by iterating over the valid i indices for X, iterating over the j indices in the intersection of X[i,_] and $\theta$, and accumulating their product in a dense vector over the i indices. The second query runs over this output and applies the sigmoid function, returning the result as a dense vector.
4.3. Execution
Once Galley has generated a physical plan, the execution is very simple. For each physical query, it first translates the expression into an STC kernel definition and calls the STC to compile it. Then, Galley injects the tensors referenced by inputs and aliases and executes the kernel, storing the resulting tensor in a dictionary by name. After all queries have been computed, we return the tensors requested in the input program by looking them up in this dictionary. In Sec. 8, we describe additional optimizations that improve this workflow by doing just-in-time physical optimization and common sub-expression elimination.
5. Logical Optimizer
Given the plan dialects above, we now describe the logical optimizer, which takes in an input program and outputs a semantically equivalent logical plan. Specifically, the logical optimizer converts each query in the input program to a sequence of logical queries where the last one produces the same output as the input query. There are many valid plans, and the optimizer searches this space to identify the cheapest one. In this section, we briefly define ”cheapest” in this context before outlining the complex space of valid logical plans that are considered. Lastly, we explain the algorithms that we use to perform this search.
5.1. Canonicalization & Pointwise Distributivity
The first step in our logical optimization is to canonicalize the input program with a few simple rules that we apply exhaustively; 1) merging nested Map operators 2) merging nested Agg operators 3) lifting Agg operators above Map operators when possible 4) renaming indices to ensure uniqueness. This compresses the input program and makes our reasoning simpler later by ensuring that operator boundaries are semantically meaningful.
Next, we consider whether to distribute point-wise expressions. This may or may not result in a better plan because it both makes operations more sparse and produces a larger expression.
Example 0.
Consider the following expression and its distributed form,
$\sum_{ij}(X_{ij}-U_{i}V_{j})^{2}=\sum_{ij}X_{ij}^{2}-2\sum_{ij}X_{ij}U_{i}V_{j}+\sum_{i}U_{i}^{2}\sum_{j}V_{j}^{2}$
If all inputs are dense, the non-distributed form is more efficient because it results in fewer terms and has the same computational cost per term. However, if $X_{ij}$ is sparse and $U_{i},V_{j}$ are dense, then the distributed form could be more efficient because all terms can be computed in time linear w.r.t. the sparsity of $X_{ij}$. Note that the squaring operation here is a point-wise function not a matrix multiplication.
Because apply pointwise distributivity can produce asymptotic performance improvements, it cannot be ignored. To address this, we perform logical optimization for both the distributed expression and the original and use the cheaper one.
5.2. Cost Model
Overall, Galley’s logical optimizer attempts to minimize the time required to execute the logical program. Because logical queries do not correspond to concrete implementations, our logical cost model aims to approximate this time without reference to the particular implementation that the physical optimizer will eventually decide on. This approximation considers on two factors: (1) the number of non-fill entries in the output tensor and (2) the amount of computation (i.e. the number of FLOPs) needed to produce the output. The former corresponds to the size of the tensor represented by Agg, $nnz(\texttt{Agg})$, and the latter corresponds to the size of the tensor represented by the MapExpr within, $nnz(\texttt{MapExpr})$. We assume that the inputs are in memory, hence there is no cost for reading inputs from disk. We then performed a simple regression to associate each of these costs with a constant, and we add them to produce our overall cost, $c$,
$cost\approx a*nnz(\texttt{Agg})+b*nnz(\texttt{MapExpr})$
To estimate $nnz(\texttt{Agg})$ and $nnz(\texttt{MapExpr})$, we introduce a sparsity estimation framework in Sec. 7.
5.3. Variable Elimination
The core of our logical optimizer is the variable elimination (VE) framework. With this view, a logical plan for an input query is defined by an order on the indices being aggregated over, i.e. an elimination order. Given this order, we can construct a valid logical plan in the following way. Iterating through the elimination order one index at a time, we (1) identify the minimal sub-query needed to aggregate it out of the expression, (2) create a new logical query representing the result of that sub-query, and (3) replace it in the original query with an alias to the result. At the end of this process, the remaining query no longer requires any aggregation and therefore is itself a logical query.
Example 0.
Consider optimizing the following matrix chain multiplication,
$E_{im}=\sum_{jkl}A_{ij}B_{jk}C_{kl}D_{lm}$
The elimination order $jkl$ corresponds to a left-to-right multiplication strategy because eliminating $j$ from the expression first requires performing the matrix multiplication between $A$ and $B$. Eliminating $k$ then requires multiplying that intermediate result with $C$, and so on. Concretely, this produces the following sequence of logical queries,
    Query(I1, Agg(+, j, Map(*, A[i,j], B[j,k])))
    Query(I2, Agg(+, k, Map(*, I1[i,k], C[k,l])))
    Query(E, Agg(+, l, Map(*, I2[i,l], D[l,m])))
Similarly, the elimination order $lkj$ corresponds to a right-to-left strategy, and the order $klj$ to a middle-first strategy.
Unlike traditional VE over sum-product queries, we need to handle complex trees of pointwise operators and aggregates. This makes the identification of minimal sub-queries challenging as we need to carefully examine the expression’s algebraic properties. Given a strategy for this, the core problem of optimizing VE is to search the space of elimination orders for the most efficient one. In the worst case, this takes exponential time w.r.t. the number of indices being aggregated over. In the following sections, we describe 1) how we identify minimal sub-queries and 2) our search algorithm for finding the optimal elimination order.
5.4. Identifying Minimal Sub-Queries
In this section, we show how to identify the minimal sub-queries (MSQ) needed to aggregate over an index. In sum-product expressions, the MSQ is just the sum over that index wrapping the product of the tensors that are indexed by it. However, it’s not immediately clear how to do this for more complex input programs which include an arbitrary mix of operators and aggregates. Fortunately, we show that identifying the MSQ corresponds to a careful traversal down the annotated expression tree, examining the algebraic properties of the operation at each node to determine how to proceed.
Annotated Expression Tree The annotated expression tree (AET) is constructed by examining the nested structure of Agg, Map, Input, and Alias nodes in the input query. We start by removing all Agg nodes and annotating their inner expressions with (Idx, Op) for each index being aggregated over. We then replace all Map nodes with their operator to get the final tree where every internal node is a function and every leaf is either an Input or an Alias.
Figure 5. Annotated expression tree for logistic regression, $\sigma(\sum_{jpc}(S_{ipc}P_{pj}\theta_{j}+S_{ipc}C_{cj}\theta_{j})$
Example 0.
Fig. 5 shows the annotated expression tree for our logistic regression example after distributing the multiplication as in Eq. (2). The sigmoid function is the outermost layer of the expression, so it appears at the top of the tree. The summations all occur just inside the sigmoid function, so they annotate the addition operator, as denoted by the red arrow. Then, each multiplication has three children which are each indexed input tensors.
Given the AET, we can identify the minimal sub-queries for a particular index by starting at the node where it’s annotated and traversing downwards according to the algebraic properties of each internal node. We now describe the traversal rules for functions which are distributive, non-distributive, and commutative with respect to the aggregation operator.
Distributive Functions When a function which distributes over the aggregate is reached (e.g. multiplication which distributes over summation), we examine how many of the children contain the current index. Each child is a subtree of the AET, and it contains the index iff it contains a tensor that is indexed by it. If one child contains the index, we simply traverse down that child’s branch, i.e. we factor the other children out of the aggregate. When multiple children contain the index, we can’t go any further down the tree, so we wrap the sub-tree rooted at that node in an aggregate and return it as our MSQ. If the function is commutative and associative, we are slightly more precise and only include the children which contain the index.
Commutative, Identical Functions When the node’s function is the same as the aggregate function and is commutative, we can push the aggregate down to each child independently. For example, if we have the expression $\sum_{i}A_{i}+B_{i}$, we can transform it into $\sum_{i}A_{i}+\sum_{i}B_{i}$. For all children which contain the index, we add the result of traversing down its branch to the MSQ and replace it with an alias to the result. If a child doesn’t contain the index, then we need to account for the repeated application of the aggregate function. To do this, we identify the function $g(x,n)=f(x,...,x)$ which represents the repeated application of our aggregate function $f$. When $f$ is addition, $g$ is multiplication. When $f$ is idempotent, $g(x,n)=x$. We then wrap each non-index child in a Map with function $g$ and the size of the index’s dimension as the second child.
Blocking Functions A function which doesn’t distribute or commute with our aggregate function is called a blocking function. When we reach a blocking function in our traversal, we simply wrap it in our aggregate and return the sub-tree as our MSQ. For example, the expression $\sum_{j}\sqrt{A_{ij}B_{jk}}$ cannot be rewritten as $\sqrt{\sum_{j}A_{ij}\sum_{j}B_{jk}}$ because $\sqrt{}$ is a blocking function.
Discussion Galley builds on and extends the FAQ framework for optimizing conjunctive queries with aggregation(Khamis et al., 2016). The FAQ framework explored the optimization of queries with the following form where each $\bigoplus^{(i)}$ is either equal to or forms a semi-ring with $\otimes$,
$\bigoplus_{v_{1}}^{\,\,\quad(1)}\cdots\bigoplus^{\,\,\quad(k)}_{v_{k}}F^{1}_{V_{1}}\otimes\cdots\otimes F^{k}_{V_{k}}$
While this captures many important problems, it lacks the flexibility needed to support a general tensor processing system. For example, consider a slightly modified version of the SDDMM kernel,
$\sum_{j}A_{ik}(B_{ij}+C+{jk})$
This expression is not an FAQ query because it mixes addition and multiplication within the pointwise expression. Similarly, our logistic regression example, $\sigma(\sum_{j}X_{ij}\theta_{j})$ cannot be expressed as an FAQ because the aggregate occurs within the $\sigma$. Galley’s logical optimizer extends this framework by accommodating arbitrary pointwise function composition and arbitrary placement of aggregates within expressions.
5.5. Restricted Elimination Orders
Lastly, before we can search for the optimal elimination order, we need to define the space of valid elimination orders. Depending on the structure of the input, the order in which indices can be eliminated might be restricted. This is due to two issues: (1) non-commutative aggregates and (2) aggregate placement. The former is when an aggregate wraps another aggregate which it doesn’t commute with. For example, if the expression is $\max_{i}\sum_{j}A_{ij}$, we have to perform the summation before handling the maximum because max and $\sum$ do not commute. The latter issue arises when an aggregate wraps another aggregate but cannot reach it via the traversal described above, e.g. $\sum_{i}\sqrt{\sum_{j}A_{ij}}$. In this case, the inner aggregate must be performed first. Collectively, these restrictions form a partial ordering on the index variables which needs to be respected when we enumerate elimination orders.
5.6. Search Algorithm
We have now simplified the complicated issue of high level optimization to the discrete problem of choosing an optimal order on the aggregated index variables. So, in this section, we present two algorithms for searching for that optimal order using the tools described above.
Greedy The greedy approach simply chooses the cheapest index to aggregate at each point. It does this by finding the minimal sub-query for each index and computing its cost. For the cheapest index, it’s minimal sub-query is removed from the expression and appended to the logical plan, then it is replaced in the remaining query with an alias to the result. This continues until no aggregates remain in the expression.
Branch-and-Bound The branch-and-bound approach computes the optimal variable order and is broken into two steps. The first step uses the greedy algorithm to produce an upper bound on the cost of the overall plan. The second step performs a dynamic programming algorithm. In this algorithm, the keys of the DP table are unordered sets of indices, and the values are a tuple of ordered lists of indices, remaining queries, and costs. We initialize the table with the empty set and a cost of zero. At each step, we iterate through the entries of the DP table and attempt to aggregate out another index. We use the cost bound from the first step to prune entries from the DP table whose cost is greater than the bound which is valid because costs monotonically increase as more indices are added to the set. At the end, we return the index order associated with the full set of indices.
6. Physical Optimizer
Each query in the logical dialect roughly corresponds to a single loop nest and materialized intermediate. However, there are still several decisions which need to be made about the way the kernel is computed; (1) the loop order over the indices, (2) the format (i.e. layout) of the result, and (3) the protocol (i.e. search algorithm) for accessing each index of each input. The physical optimizer makes these decisions.
6.1. Loop Order
The loop order of a tensor kernel determines which order the inputs are accessed in. A good order will result in early pruning of iterations due to early intersection of sparse inputs. Intuitively, this is similar to selecting a variable order for a worst-case optimal join algorithm.
Cost Model
The cost of a loop order is equal to how many iterations each level of the loop nest incurs.
Example 0.
Consider matrix chain multiplication over three sparse matrices, $A$, $B$, and $C$.
(3) $D[il]=\sum_{jk}A[ij]*B[jk]*C[kl]$
Suppose that $A$ has only a single non-zero entry and that $B$ and $C$ have 5 non-zero entries per column and per row. Then, the loop order $ijkl$ is significantly more efficient than $lkji$. In the former, the first two loops, over $i$ and $j$, only incur a single iteration because they are bounded by the size of $A$. The third and fourth occur $5$ and $5^{2}$ times respectively because there are only $5$ non-zero $k$’s per $j$ in $B$ and $5$ non-zero $l$’s per $k$ in $C$. In the latter, the first two loops iterate over the full matrix $C$ despite most of those iterations not leading to useful computation.
The last piece of our cost model is the cost of transposition. If an input’s index order is not concordant with the loop order, it must be transposed before the query can be executed. This imposes a cost that is linear in that input’s size.
Optimization Algorithm
To optimize the loop order, we combine this cost model with a branch-and-bound, dynamic programming algorithm. In the first pass, it selects the cheapest loop index at each step until reaching a full loop order. This produces an upper bound on the optimal execution cost which we use to prune loop orders in the second step. This step applies a dynamic programming algorithm. Taking inspiration from Selinger’s algorithm for join ordering, each key in the DP table is a set of index variables and a set of inputs. The former represent the loops that have been iterated so far, and the latter represents the inputs which will need to be transposed based on the optimal prefix.
6.2. Intermediate Formats
Once the loop order has been determined, we select the optimal format for each query’s output. Because we operate within the fiber tree abstraction, this means that we select a format for each index of the output (e.g. dense vector, hash table, etc.). There are two factors which affect this decision: (1) the kind of writes being performed (sequential vs random) (2) the sparsity of the resulting tensor. The former is important because many formats (e.g. sorted list formats) only allow for sequential construction. This means that they can only be applied if the indices of the output form a prefix of the loop order. The latter factor balances the fact that denser formats tend to be more efficient due to their memory locality and simplicity while sparser formats are asymptotically better when dealing with significantly sparse outputs. To describe this trade-off, we hand selected sparsity cutoffs between fully sparse, bytemap, and fully dense formats. To determine a particular output index’s format, we first determine the sparsity at this index level and use our cutoffs to determine which category of formats to consider. Then, we check whether we are performing sequential or random writes and select the most efficient format that supports our write pattern.
6.3. Merge Algorithms
The final decision for the physical optimizer to make is the algorithm it will use to perform each loop’s intersection. While there are some more complex strategies, we adopt a minimal approach and select a single input to iterate over for each loop. The remaining inputs are then probed into. We make this selection by estimating the number of non-zero indices that each input has, conditioned on the indices in the outer loops. This is similar to the approach taken in (Wang et al., 2024) for optimizing WCOJ.
7. Sparsity Estimation
In this section, we describe how our system performs the sparsity estimation that guides our logical and physical optimizers. We start by discussing the subtle correspondence between sparsity estimation and cardinality estimation. Then, we describe a minimal interface for sparsity estimation inspired by this correspondence. Lastly, we discuss two implementation of this framework; the uniform estimator and the chain bound.
7.1. Sparsity & Cardinality Estimation
Sparsity estimation is highly related to cardinality estimation in databases. However, translating methods for the latter to the former requires analyzing the algebraic properties of our tensor programs. For example, let $A_{ij}$ and $B_{jk}$ be sparse matrices with a fill value of 00 and $R_{A}(I,J)$ and $R_{B}(J,K)$ be relations which store the indices of their non-zero entries. Suppose we’re performing matrix multiplication,
$C_{ijk}=A_{ij}B_{jk}$
Then, the number of non-zero values in $C$ is precisely equal to the size of the following conjunctive query,
$nnz(C)=|R_{A}(I,J)\bowtie R_{B}(J,K)|$
The correspondence is due to the fact that 00 is the annhilator of multiplication (i.e. $x*0=0\forall x$), so any non-zero entry $ijk$ in the output must correspond to a non-zero $ij$ in $A$ and a non-zero $jk$ in $B$. Now, consider the following instead,
$C_{ijk}=A_{ij}+B_{jk}$
In this case, a nonzero $ijk$ in the output can result from a non-zero $ij$ in $A$ or a non-zero $jk$ in $B$. In traditional relational algebra where relations are over infinite domains, this kind of disjunction would result in an infinite relation. However, tensors have finite dimensions, so we can introduce relations that represent the finite domains of each index, e.g. $D_{i}=\{1,...,n_{i}\}$. This allows us to represent the index relation of the output as,
$nnz(C)=|(R_{A}(I,J)\bowtie D_{k}(K))\cup(D_{i}(I)\bowtie R_{B}(J,K))|$
Lastly, we can translate aggregations to the tensor setting as projection operations. If we have the following statement,
$C_{ik}=\sum_{j}A_{ijk}$
We can express the non-zeros entries of $C$ as,
$nnz(C)=|\pi_{I,K}(R_{A}(I,J,K))|$
7.2. The Sparsity Statistics Interface
We use our statistics interface to annotate an expression with stats objects at every node of the AST in a bottom-up fashion. Each stats object then represents the sparsity patterns of the intermediate tensor output from that node. Surprisingly, to support sparsity estimation over the varied workloads and arbitrary operators of tensor algebra, we only need to implement a few core functions; 1) A constructor which produces statistics from a materialized tensor. This produces statistics for Input and Alias nodes. 2) A function for annihilating Map nodes (i.e. those whose children’s fill values are the annihilator of its pointwise function) which merges the children’s statistics. 3) A function for non-annihilating Map nodes which merges the children’s statistics. 4) A function for Agg which adjusts the input’s statistics to reflect an aggregation over some set of indices. 5) An estimation procedure that estimates the number of non-fill entries based on statistics about a tensor.
7.3. Supported Sparsity Estimators
7.3.1. Uniform Estimator
The simplest statistic that can be kept about a tensor is the number of non-fill (e.g. non-zero) entries. The uniform estimator uses only this statistic and relies on the assumption that these entries are uniformly distributed across the dimension space. This corresponds to System-R’s cardinality estimator with the added assumption that the size of the index attribute’s active domain is equal to the size of the dimension (Selinger et al., 1979).
Constructor
Given a tensor $A_{i_{1},\ldots,i_{k}}\in\mathbf{R}^{n_{i_{1}}\times\ldots\times n_{i_{k}}}$, we simply count the non-fill values in the tensor, $nnz(A)$, and note the dimension sizes $n_{i_{1}},\ldots,n_{i_{k}}$.
Map (Annihilating)
To handle an annihilating pointwise operation, we calculate the probability that a point in the output was non-fill in all inputs, then multiply this with the dimension space of the output. For a set of inputs $A^{(1)}_{I_{1}}\ldots A^{(l)}_{I_{l}}$ and output $C_{I_{C}}$ where each $I_{j}$ is a set of indices, this probability is
$nnz(C)\approx\left(\prod_{i\in I_{C}}n_{i}\right)\cdot\left(\prod_{j}\frac{nnz(A_{j})}{\prod_{i\in I_{j}}n_{i}}\right)$
Map (Non-Annihilating)
To handle an non-annihilating pointwise operation, we calculate the probability that an entry in the output was fill in all inputs. Then, we take the compliment to get the probability that it was non-fill in all inputs and multiply this with the output dimension space. Using the notation from above,
$nnz(C)\approx\left(\prod_{i\in I_{C}}n_{i}\right)\cdot\left(1-\prod_{j}\left(1-\frac{nnz(A_{j})}{\prod_{i\in I_{j}}n_{i}}\right)\right)$
Aggregate
Given an input tensor $A_{I}$ which we are aggregating over the indices $I^{\prime}$, we compute the probability that an output entry is non-fill by calculating the probability that at least one entry in the subspace of the input tensor wasn’t fill.
$nnz(C)\approx\left(\prod_{i\in I\setminus I^{\prime}}n_{i}\right)\cdot\left(1-\left(1-\frac{nnz(A_{I})}{\prod_{i\in I}n_{i}}\right)^{\prod_{i\in I^{\prime}}n_{i}}\right)$
Estimate
The estimation function simply returns the cardinality statistic stored about the current tensor.
7.3.2. Degree Statistics & The Chain Bound
We keep degree statistics as the default in Galley, and we use them to compute upper bounds on the number of non-fill entries in intermediate expressions. A degree statistic, denoted $D_{A}(X|Y)$, stores the maximum number of non-fill entries in the $X$ dimensions conditioned on the $Y$ dimensions for a tensor $A$. For example, if you have a matrix $A_{ij}$, then $D_{A}(i|j)$ is the maximum number of non-fill entries per column, and $D_{A}(ij|\emptyset)$ is the total number of non-fill entries in the matrix. This approach follows work in cardinality bounding which has been shown to produce efficient query plans in the relational setting (Deeds et al., 2023; Hertzschuch et al., 2021; Cai et al., 2019).
Constructor
We first compute the boolean tensor representing the input’s sparsity pattern. Then, to calculate each degree statistic, we sum over the $X$ dimensions and take the maximum over the $Y$ dimensions. The set of degree statistics for a tensor $A_{I}$ is denoted $\mathcal{D}_{A_{I}}$
Map (Annihilating)
Annihilating map operations function as conjunctive queries with respect to the sparsity patterns of the inputs. Therefore, any degree statistics that are valid for an input are also valid about the output, so we compute the statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by a union,
$\mathcal{D}_{C}=\bigcup_{j}A^{(j)}_{I_{j}}$
Map (Non-Annihilating)
In this case, we need to be more careful to ensure that we maintain our upper bounds. First, we extend the degree constraints from each input to cover the full set of indices. For example, if we have $DC_{X}(i|j)$ and want to extend it to the dimension $k$, then we compute $D_{X}(ik|j)=DC(i|j)*n_{k}$. Then, we compute degree statistics about the output, $C$, from the inputs $A^{(1)}_{I_{1}},\ldots,A^{(k)}_{I_{k}}$ by addition,
$D_{C}(X|Y)=\sum_{j}D_{A^{(j)}}(X|Y)$
Estimator
We calculate our upper bound (eq. perform sparsity estimation) using the breadth-first search approach described in (Chen et al., 2023). Intuitively, each set of indices forms a node in the graph, and each degree constraint is a weighted edge from $Y$ to $X$. Our search begins with the empty set, then we use breadth first search to find the shortest weighted path to the full set of indices $I$. The product of the weights along this path bound the number of non-zeros in the result.
8. Additional Optimizations
In this section, we outline a few additional techniques we apply to improve the optimization and execution of tensor programs.
8.1. Just-In-Time Physical Optimization
When optimizing a large program, estimates of sparsity become less reliable because they require more steps of inference on top of grounded inputs. This poses a challenge for the physical optimizer as it estimates the cost of different loop orders, formats, and access protocols. To ameliorate this, Galley performs a very simple form of adaptive optimization by waiting to perform physical optimization on each logical query until all of its aliases have already been executed. This allows it to calculate the size of prior intermediate results and update the relevant statistics before performing physical optimization. This makes that cost estimates in that optimization more accurate and produces a better physical plan.
8.2. Common Sub-Expression Elimination
We take a straightforward approach to avoiding redundant computation. Before executing each physical query, we first canonicalize the right hand side, hash it, and compute its hash value. We use this to check a cache to see if it has been executed already. If so, we immediately return the value from the cache rather than re-executing the query. Similarly, after computing a query, we store the result in this cache for later reuse.
Figure 6. Machine Learning Algorithms Over Joins
9. Experimental Evaluation
In this section, we evaluate the effectiveness of our optimizer on a variety of workloads: (1) machine learning algorithms, (2) subgraph counting, and (3) breadth-first search. Compared to hand-optimized solutions and alternative approaches, Galley is highly computationally efficient while only requiring a concise, declarative input program from the user. Overall, we show that Galley:
[…]
  • Selects optimal tensor formats for intermediates, beating both the dense and sparse formats for 3/5 graphs
Experimental Setup
These experiments are run on a server with an AMD EPYC 7443P Processor and 256 GB of memory. We implemented Galley in the programming language Julia, and the code is available at https://github.com/kylebd99/Galley. We used the sparse tensor compiler Finch for execution, and all methods are executed using a single thread. Unless otherwise stated, Galley uses the chain bound described in Sec. 7.3.2 for sparsity estimation. Experiments for all methods are run three times and the minimum execution time is reported. This removes the compilation overhead which we address separately in Fig. 9.
9.1. Machine Learning Algorithms
To explore end-to-end program optimization, we experiment with simple machine learning algorithms over joins, represented entirely in tensor algebra. For this, we use the TPC-H benchmark and consider two join queries, a star query and a self-join query. First, we perform a star join over the line items table to bring together features about suppliers, parts, orders, and customers. This is expressed as follows where $L,S,P,O,$ and $C$ are tensors representing the lineitems, suppliers, parts, orders, and customers tables, respectively.
$X_{ij}=\sum_{spoc}L_{ispoc}(S_{sj}+P_{pj}+O_{oj}+C_{cj})$
The non-zero values in $S,P,O$ and $C$ are disjoint along the $j$ axis, so the addition in this expression serves to concatenate features from each source, resulting in 139 features after one-hot encoding categorical features. The self-join query compares line items for the same part based on part and supplier features. In this case, the feature data is a 3d tensor because the data points are keyed by pairs of line items.
$X_{i_{1}i_{2}j}=\sum_{s_{1}s_{2}p}L_{i_{1}s_{1}p}L_{i_{2}s_{2}p}(S_{s_{1}j}+S_{s_{2}j}+P_{pj})$
Given these definitions, we consider a range of ML algorithms; 1) linear regression inference 2) logistic regression inference 3) covariance matrix calculation and 4) neural network inference. As a comparison point, we also implement two versions of each of these using the Finch compiler. The dense version uses a fully dense matrix to represent the feature matrix whereas the sparse version uses a sparse level for the features to compress the one-hot encoded features. Fig. 6 shows that Galley is 2100×2-100$\times$2 - 100 × faster for all but one experiment. This performance improvement is derived from Galley pushing the computation into the definition of $X$. For example, linear regression inference is fundamentally a matrix-vector multiplication, and Galley pushes the vector multiplication all the way down to the feature matrices $S,P,O,$ and $C$. Note that we do not provide a dense implementation for the self-join query because it runs out of memory due to the highly sparse $i_{1}$, $i_{2}$ dimensions.
Figure 7. Subgraph Counting Execution Time
Figure 8. Subgraph Counting Optimization Time
Figure 9. Subgraph Counting Compilation Time
Figure 10. BFS Runtime
9.2. Subgraph Counting
To test Galley’s performance on large, sparse problems, we implement a few common sub-graph counting benchmarks. The conversion from sub-graph counting to sparse tensor algebra is straightforward. Suppose you are counting the occurrences of $H(V,E)$ in a data graph $G$ with adjacency matrix $M$, then we can represent the count as,
$c=\sum_{v_{i}\in V}\prod_{(v_{i},v_{j})\in E}M_{v_{i}v_{j}}$
To handle vertex labels, we add sparse binary vector factors for each labeled vertex. We use subgraph workloads from the G-Care benchmark and the paper ”In-Memory Subgraph: an In-Depth Study”(Sun and Luo, 2020; Park et al., 2020). We restrict the latter benchmark to query graphs with up to 8 vertices, hence the ”_lite” suffix . Because this is a relational workload, we compare with DuckDB, a modern OLAP database (Raasveldt and Mühleisen, 2019). To tease apart the impact of logical optimization from physical optimization and our use of Finch, we provide a version of Galley that executes each logical query with a SQL query run on DuckDB. Lastly, we provide results for the greedy logical optimizer as well.
In Fig. 7, we see median execution times which are up to 100×100$\times$100 × lower when comparing Galley to DuckDB. When using Galley’s physical optimizer with Finch as opposed to using DuckDB as an execution engine, we still see up to 10×10$\times$10 × lower median execution time. The former shows the benefit of Galley’s logical optimizer and the variable elimination framework. The latter demonstrates the benefits of using STCs which enable a wide range of formats and produce highly efficient code. Further, DuckDB hits the 300 second timeout on 196 out of 400 queries in the youtube_lite benchmark. This is reduced to 5 and 18 timeouts when using Galley with a DuckDB and Finch execution engine, respectively. The greedy optimizer performs similarly to the exact optimizer on these workloads with the latter providing improvements on the slowest queries. Fig. 8 shows the mean optimization time for each of the methods on each of the workloads. Galley has a mean optimization time of less than .15.15.15.15 seconds across all workloads, approaching the time taken by the highly efficient DuckDB optimizer.
Lastly, because it performs compilation using Finch at runtime, Galley incurs a compilation latency when it first requests each tensor kernel implementation. Fortunately, these kernels are cached automatically by Finch, reducing this cost when workloads repeatedly use similar kernels. We show the mean compilation time for each subgraph workload in Fig. 9. On the simpler workloads which often reuse kernels, this cost is minimal. However, the more complex workloads both reuse kernels less and require compiling more complex kernels, significantly increasing compilation time. This suggests a need for faster STC compilation or the use of an interpreted engine for complex queries.
9.3. Breadth First Search
To demonstrate the importance of selecting optimal formats, we implement a simple push-based breadth first search algorithm using Galley and different hand coded Finch implementations. Both systems are provided with a single iteration at a time and the total execution time across all iterations is reported. This means that the core optimization question is how to represent the vector of visited vertices and the vector of frontier vertices. The former grows monotonically over the course of the algorithm while the number of non-zeros in the latter forms a curve, peaking in the middle iterations. We provide two implementations of Finch 1) using a sparse vector for both intermediates and 2) using a dense vector for both intermediates. Fig. 10 shows that Galley is faster than both dense and sparse approaches for 3 of the 5 graphs and is competitive for all graphs. Note that this includes the optimization time for Galley.
10. Conclusion & Limitations
In this paper, we presented Galley, our system for declarative sparse tensor programming. We described and then demonstrated how it optimizes high-level program structure with its logical optimizer and how it lowers that program to an efficient implementation with its physical optimizer. These decisions are guided by sparsity estimates of intermediate expressions, and we show that these estimates can be computed for arbitrary tensor algebra programs by implementing a minimal 5-function interface. Lastly, we present an experimental evaluation of this work on ML algorithms over structured feature data, sub-graph counting, and breadth-first search.
There are a few optimizations which we are excited to bring to Galley in the future. Currently, it does not support complex loop structures (e.g. a single outer FOR loop which wraps multiple inner FOR loops) or parallelism. Both of these areas can benefit from cost-based optimization, and we’re excited to explore them in future work. Similarly, we do not consider memory constraints during the optimization process, but our focus on cardinality bounds provides an exciting avenue for approaching this in the future.